### v2

# VGA: Vision GUI Assistant - Minimizing Hallucinations through Image-Centric Fine-Tuning

Abstract
Recent advances in Large Vision-Language Models (LVLMs) have significantly improve performance in image comprehension tasks, such as formatted charts and rich-content images. Yet, Graphical User Interface (GUI) pose a greater challenge due to their structured format and detailed textual information. Existing LVLMs often overly depend on internal knowledge and neglect image content, resulting in hallucinations and incorrect responses in GUI comprehension. To address these issues, we introduce VGA, a fine-tuned model designed for comprehensive GUI understanding. Our model aims to enhance the interpretation of visual data of GUI and reduce hallucinations. We first construct a Vision Question Answering (VQA) dataset of 63.8k high-quality examples with our propose Referent Method, which ensures the model’s responses are highly depend on visual content within the image. We then design a two-stage fine-tuning method called Foundation and Advanced Comprehension (FAC), to enhance both the model’s ability to extract information from image content and alignment with human intent. Experiments show that our approach enhances the model’s ability to extract information from images and achieves state-of-the-art results in GUI understanding tasks. Our dataset and fine-tuning script will be released soon.
VGA: Vision GUI Assistant - Minimizing Hallucinations through Image-Centric Fine-Tuning
Figure 1: Data generation Method
1 Introduction
Large Vision-Language Models (LVLMs) have recently emerged as a powerful approach for various multimodal tasks. These models acquire textual knowledge through pre-training and develop image understanding abilities during instruction-tuning Cao et al. (2023); Liu et al. (2023). LVLMs can effectively process both visual and linguistic information by using a visual-language projector, which maps different types of data into a shared latent space. For instance, LLaVA Liu et al. (2024a) and InstructBLIP Dai et al. (2023) have demonstrated strong zero-shot capabilities in tasks like image captioning, visual reasoning, and complex conversations. UniChart Masry et al. (2023), on the other hand, has demonstrated the powerful capabilities of LVLMs in the understanding of formatted charts.
With the proliferation of mobile applications, the importance of Graphical User Interfaces (GUIs), which serve as a critical bridge between end users and applications, has increasingly garnered scholarly attention. GUIs, characterized by their structured layouts, rich graphical and textual content, and the inclusion of human operational logic Banerjee et al. (2013), present a complex challenge: Can the success of LVLMs be applied to the GUI domain?
Existing GUI comprehension method has transferred from code-centric base to visual base. While code-centric base approaches offer a foundational outline Chen et al. (2020b), they fail to fully capture the essence of GUI content, making visual-based methods more advantageous Yu et al. (2020)Bajammal et al. (2022). Compared to code-centric methods, visual-based approaches can convey the actual user interface interactions and mirror the user’s direct experience with the GUI. For instance, LabelDroid Chen et al. (2020a) utilizes deep learning to predict labels for image-based buttons from a variety of commercial apps available on Google Play. Similarly, TANGO Cooper et al. (2021) employs custom computer vision (CV) and text retrieval techniques to analyze visual and textual information on mobile screens.
However, traditional visual GUI comprehension focuses on identifying GUI components and fails to comprehend the graphical and textual information, layout, and interaction context within the interface. True comprehension demands that the model interpret the GUI in a manner akin to human perception. Existing LVLMs demonstrate significant advantages in this domain as they can understand both formatted charts and rich textual content while leveraging natural language aligned with human interpretations Zhang et al. (2024). Nevertheless, directly applying conventional LVLMs to GUI comprehension tasks often results in hallucinations, where the model over-relies on its pre-trained knowledge and neglects the actual visual content, leading to inaccurate or irrelevant responses Shahgir et al. (2024); Zhang et al. (2023).
To address these issues, we propose VGA, a model fine-tuned on a self-construct 63.8K dataset, using a novel training method we design. During the dataset construction, we employ knowledge distillation from large language models to construct 63.8k dataset based on Rico Deka et al. (2017), and we adopt the Referent Method to enhance the model’s focus on image content by using visual and position information in constructing dataset, thereby increasing the relevance and accuracy of the responses. In the training process, we apply a Foundation and Advanced Comprehension (FAC) approach: the Foundation Stage enhances the model’s understanding of GUI image, while the Advanced Comprehension stage improves the model’s ability to respond to complex questions based on its understanding of the GUI. Additionally, we employ a task progression and reinforcement approach to create logical chains between tasks and within responses to strengthens the model’s ability to reason and infer relationships within the provided context. Through these integrated strategies, VGA achieves state-of-the-art performance in GUI comprehension tasks. Our primary contributions are as follows:
  • Large-scale GUI corpus for LVLM Fine-tuning. To fine-tune our model, we introduce a large-scale GUI corpus that includes a diverse array of apps accompanied by corresponding text descriptions and dialogues.
  • A Fine-tuned LVLM for GUI task. We propose VGA, a LVLM for GUI comprehension, fine-tuned to fulfill both granular low-level and strategic high-level goals specialized for graphical user interfaces.
  • A Fine-tune Method to efficient improve LVLMs performance We propose a two stage fine-tune method based on our dataset to achieve better understand of images.
  • Performance enhancements on real-world GUI tasks. As shown in experiment  3, we apply our VGA to a GUI comprehension bench, yielding promising results.
2 Related Works
2.1 GUI Comprehension
GUI comprehension is essential in mobile agent design to ensure app quality and user experience. It focuses on assessing interface components such as buttons, text boxes, and menus to verify their operational performance, visual design, and usability, enabling mobile agents to interact effectively with the app interface Yu et al. (2023b). The software engineering community has been long-term focused on the improvement of mobile app GUI comprehension effectiveness in all aspects since last century Memon et al. (1999). The main purpose of these efforts is to understand GUI elements to advance GUI automation, including testing and other applications Arnatovich and Wang (2018); Vásquez et al. (2018); Said et al. (2020).
Existing methods for GUI comprehension can be divided into code-based approaches, which rely on functional specifications or code analysis Paiva et al. (2005); El Ariss et al. (2010), and visual-based approaches, which improve performance by considering the GUI’s visual representation through techniques like template matching and OCR Cheng et al. (2019). However, these methods still lack the ability to interact with GUIs in a human-like manner. Recently, Large Vision-Language Models (LVLMs) have emerged as a promising approach by combining visual and textual understanding. These models leverage vast prior knowledge gained from pre-train to comprehend human intentions and interact more naturally with GUI elements Wang et al. (2024); Cui et al. (2024).
In our work, we leverage the strengths of LVLMs and enhance their performance in GUI comprehension using our constructed dataset and our designed training methods.
2.2 Large Vision-Language Models
The introduction of the transformer architecture has revolutionized natural language processing, enabling models to efficiently capture long-range dependencies and contextual information. This advance laid the groundwork for pre-trained Large Language Models (LLMs). The pre-train, fine-tune, and predict paradigm Liu et al. (2021), exemplified by models such as GPT Radford et al. (2018) and BERT Devlin et al. (2018), led to significant improvements in language understanding and generation Cheng et al. (2023); Yao et al. (2023); Arefeen et al. (2024); Schick and Schütze (2020); Yu et al. (2023a). Subsequent models like GPT-4 Achiam et al. (2023), Llama Touvron et al. (2023), and Qwen Bai et al. (2023) have further expanded these capabilities. Explorations into the Mixture of Experts (MoE) architecture Shen et al. (2023); Jiang et al. (2024) continue to enhance the scalability of Transformer-based models.
Besides, using transformers to map visual information into a semantic space has ensured the robust extension of LLMs into multimodal applications Yin et al. (2023). This enables models to process vision information, exemplified by multimodal recognition models like CLIP Radford et al. (2021), which align semantic information from images with text in a unified vector space. Advanced methodologies by Sun et al. (2023) and Fang et al. (2023) have further refined these capabilities.
Large Vision-Language Models (LVLMs) harness the strengths of both LLMs and visual feature encoders, utilizing various methods to project visual data into an LLM-comprehensible space Popescu et al. (2009); Liu et al. (2023, 2024b); Alayrac et al. (2022); Wang et al. (2023). Models like Flamingo Alayrac et al. (2022) and CogVLM Wang et al. (2023) exemplify the architectures that achieve this integration. Additionally, UniChart Masry et al. (2023) excels in chart comprehension by leveraging LVLMs to accurately interpret formatted charts and complex informational structures. Similarly, LLaVA Liu et al. (2023) integrates pre-trained visual and language models to enhance the understanding of multimodal inputs, achieving superior performance in visual and textual comprehension tasks.
Inspired by these work, we investigate the performance of LVLMs in the context of complex GUI interfaces. Our study focuses on evaluating how well these models understand and interact with GUIs that feature structured layouts and rich interactions, aiming to further enhance their applicability and accuracy in real-world scenarios.
3 GUI Dataset Construction
Inst dataset Model Number Conv dataset Model Number
description_inst GPT-4 3k Conv_simple GPT-4 5.4k
bound_inst GPT-4 4.3K Conv_complex GPT-4 11K
function_inst GPT-4 2K Conv_4o_long GPT-4o 10K
testing_inst GPT-4 5K Conv_4o_short GPT-4o 10K
function_inst_4o GPT-4o 8K conv_4o_miss GPT-4o 5k
Total 63.8K
Table 1: Dataset composition
Existing general GUI dataset cannot be directly use for LVLM training. To address this gap, we propose a novel approach to generating high-quality GUI comprehension dataset that meets the urgent demand of LVLM training.
3.1 Existing General Dataset
The Rico Deka et al. (2017) dataset is a huge GUI dataset created to support research in mobile app design and development, including areas such as GUI design, interaction. It consist of 66k unique GUI screens and 3M elements from 27 categories, over 9.3k applications. Each GUI comprise a screenshot and an augmented Android view hierarchy that capture all of the elements comprising a GUI, their properties, and relationships between them.
However, the rich data provided by Rico cannot be directly utilized to train LVLMs as the text data in Rico do not align with human perception in the same manner as VQA datasets. This necessitates the transformation of this data into a format compatible with VQA dataset.
3.2 Propesed GUI VQA Dataset Generation
Inspired by the works like LLaVA Chen et al. (2024) and ShareGPT4V Chen et al. (2023), we use state-of-the-art models like GPT-4OpenAI (2024) to generate training data. More details about this approach can be found in Appendix A.1
In constructing high quality VQA datasets for GUI comprehension tasks, multiple quality dimensions must be considered:
Interaction: The ability to understand user interactions with GUI elements is crucial for the model to comprehend GUI functionality.
Positional: Accurate description of element locations within the interface is crucial, as relative position imply connections between GUI elements, such as buttons next to text boxes indicating related actions.
Visual Detail: Accurate representation of visual information, such as shape and color, is crucial. This allows the model to distinguish between elements and focus on their design attributes within the GUI.
Overall Function: Comprehensive overall functional descriptions enable the model to grasp the intended purpose and usage patterns of the interface, facilitating more accurate and effective comprehension.
The Android view hierarchy in the Rico dataset provides detailed descriptions of GUI elements, allowing us to obtain precise positional data and interaction details. Therefore, we use GPT-4 (text-only) to extract this textual information and GPT-4o to extract visual information as shown in Figure 1. Our approaches are as follows:
  • Data Pre-processing: We remove elements in the Android View Hierarchy that have the ‘visible-to-user’ attribute set to ‘false’, and normalize the ‘bounds’ values to a range of 0 to 1. Additionally, for all elements with the ‘clickable’ attribute set to ‘true’, we add a ‘click_coordinate’ attribute, calculated as the midpoint of the range indicated by the ‘bounds’.
  • Data Generation by GPT-4 (text-only): Inspired by the method of LLaVA Liu et al. (2024b), We leverage the advanced capabilities of GPT-4 models (gpt-4-0125-preview and gpt-4-1106-preview) to generate VQA data. We employ these models to parse Android View hierarchy, extract precise element locations and user interaction information. Our approach ensures a accurate description of GUIs, including positional relationships and user interaction.
  • Data Generation by GPT-4o (vision-model): Inspired by ShareGPT4 Chen et al. (2023), we employ GPT-4o to directly process images and generate QA pairs based on the visual content. we capture the design patterns between different GUI elements.
Example
[Uncaptioned image] [Uncaptioned image]
User How can I give a like to this video? How can I login?
LLaVA-v1.6-34B Open the video you want to like. Look for a thumbs-up or heart icon, which is usually located near the bottom of the screen Open the website or app of the social media platform you want to log in to. Look for the "Login" or "Sign In" button, which is usually located at the top right corner of the page.
GPT-4V there should be a heart icon or similar button typically found below the video or in a sidebar. Tapping on this icon… Open the App or website…
Find login section…
Enter your credentials…
VGA You can click the heart-shaped icon located on the right side of the screen, just below the person’s image. This page seems to be focused on displaying content and interacting with it, but it doesn’t appear to have a login function.
Table 2: Example of hallucination in GUI comprehension, red means element in image relate to answer and orange means location of the element which indicate if model really extract information from image.
Our GUI comprehension dataset consists of 63.8k images, covering a diverse range of apps and tasks. It includes 22.3k instruction-following data pairs and 41.4k conversation data pairs. Of the instruction data, 35.8% is generated by GPT-4o, and 60.2% of the conversation data is produced by GPT-4o. Table 1 illustrates the composition of our dataset.
4 GUI Comprehension Task Design
In this section, we discuss how hallucinations arise in GUI understanding tasks and introduce our dataset design, explaining how it addresses this issue. The design of our dataset is related to our fine-tuning method. For more details, please see Section  5.
4.1 Hallucination in GUI Comprehension
The primary cause of hallucinations in LVLMs can be attributed to their dependence on response patterns learned by traditional Large Language Models (LLMs). LLMs are typically trained on massive corpora of pure text data, where they learn to generate responses based on textual context and patterns. When a visual module is introduced, these learned textual response patterns can adversely influence the model’s behavior during multimodal tasks, as shown in Table 2.
In a GUI, the structured layout conveys relationships between image and text elements, such as a number beneath a heart indicating likes. Misunderstanding these positional relationships can cause incorrect model responses. Moreover, action-indicating text often guides user behavior in GUI design, but the attention method might cause the model to over focus on similar words in the question. For instance, the model might suggest clicking a "login" button when you want to login in a page even if it doesn’t exist.
Referent Method For hallucination Reduction: To address the previously mentioned issue, we propose the referent method. GUI design often uses layout, shape, and color to distinguish elements, providing critical visual details to guide user interactions. By explicitly incorporating coordinates, shapes, colors, and relative positions between elements in the construction of our dataset, we aim to enhance the model’s focus on the image content when generating answers. Thereby reducing the chances of hallucination. During the design of our dataset, we ensured that most responses involving GUI elements include at least one of the following referents, aligning the element information with the image content:
Shape: The shape of the element, like a rectangular button with rounded corners.
Color: The color of the element, for example, a blue button with white text.
Position: The exact coordinates or bounds of the element, format as <x,y> and <x1,y1,x2,y2>.
Relative Position: The position of the element in relation to other elements, such as being below the text input field.
4.2 Foundation Task For Knowledge Injection
The Foundation Task dataset is an instruction-following dataset that fix the command and response formats, ensuring that only the image content and response information vary dynamically. This dataset enhance model to identity the linking between specific data characteristics with visual attributes and understand the functions provided by the GUI. Our foundation task can be divided as follow:
  • Simple Description: This task involves providing a basic description of the GUI’s layout, identifying the function and placement of components. This foundational task facilitates deeper analytical tasks, as all GUI-related tasks rely on a clear understanding of the overall layout.
  • Bounds & Location: This task adds complexity by requiring the model to incorporate precise descriptions of element bounds and coordination. By introducing this, our approach ensures that the model focuses on the accurate positional information within the images.
  • Overall Function: This task requires the model to grasp the individual functions of each GUI element based on overall function of the entire interface. Identical buttons might serve different purposes depending on their specific context within a GUI. This enables the model to predict the design intent and functionality of the entire interface, resulting in more contextually aware and precise responses.
4.3 Advanced Task For Human Alignment
Building upon the capabilities developed in the foundation tasks, the advanced comprehension tasks are designed to simulate real-world human interactions and complex dialogues within GUI comprehension tasks. This tasks are characterized by more flexible and contextually rich questions and responses compared to the foundational tasks. While the basic capabilities required are similar to those in the Foundation task, the advanced comprehension tasks emphasize the application and deeper understanding of these foundational skills. Our advanced comprehension tasks can be divided into two categories, distinguished by the generating models and its "referent":
  • GPT-4 (Text-Only) Generated Tasks: This category consists of conversations generated by GPT-4 (text-only), based on pre-processed Android View Hierarchy data. Leveraging the detailed information available in the Android View Hierarchy, this approach generates a variety of QA pairs that accurately reflect the textual and positional data and interactions.
  • GPT-4o (Image-Based) Generated Tasks: This category comprises dialogues generated by GPT-4o (image-based). Given that vision models often struggle with capturing precise positional information of elements, the question and response in this dataset focus on relative positions, shapes, and colors of elements.
5 Foundation and Advanced Comprehension
Existing fine-tuning methods in LVLM treat question tokens and image tokens equally, which neglects the need for LVLM to prioritize image information. As a result, models may understand tasks well but generate responses based solely on questions. This results in hallucinations in GUI comprehension (see Table 2). To address this, we propose a two-stage training method. The first stage aligns responses with image content. The second stage aligns responses with human intent. We also adopt the chain-of-thought method to enhance the model’s reasoning capability. Our approaches are as follows:
Foundation Stage: In this stage, we fine-tune the model with the Foundation task dataset, which covers essential GUI comprehension capabilities. The fixed instruction and response format (detailed in sec 4.2) guarantees that only the images and response content vary for each sample. This approach makes response generation focusing more on image content, and builds foundational skills for GUI comprehension.
Advanced Comprehensive Stage: In this stage, we introduce multi-turn dialogues and complex questions from the advanced comprehension dataset, where each sample has a flexible question and answer format aligned with human intent.. In this phase, model learns human questioning patterns and utilizes foundational skills acquired from the first stage to generate responses. By employing the Referent Method (detailed in sec 4.1), our approach ensures that the model’s answers include more specific image details, thereby reducing hallucinations and improving the accuracy of the responses.
Figure 2: Loss convergence of model trained with foundation task and without during advanced task training.
Figure 3: Loss convergence of model trained with mixed foundation task and advanced task data.
Task Progression and Reinforcement: We organize tasks in a sequence from simple to complex, ensuring that the model progressively builds the necessary foundational skills before training on more advanced tasks. In complex tasks, we start with some related simple tasks to reinforce foundational knowledge, followed by the real response which demands deeper analysis and synthesis of the information gathered from these foundational tasks.
Model First Second Third Average
GPT-4o 80.68 80.45 81.14 80.75
GPT-4V 81.82 81.14 82.50 81.82
MiniCPM 63.86 64.09 64.32 64.09
llava-next 53.41 53.18 53.86 53.48
idefics2-8b 39.77 43.86 41.81 41.82
VGA-7b-v1* 59.00 59.75 57.75 58.5
VGA-7b-v1 90.68 90.68 91.17 90.83
Table 3: Scores on GUI Bench.
5.1 Experiments Setup
We use our training method and dataset to fine-tune llava-v1.6-mistral-7b Liu et al. (2024a). Table 5 shows the hyperparameters we use during training. In Appendix A.2, we show the loss convergence behavior under various learning rates and batch sizes during training. All our experiments are carried out using one A100(80GB) machine.
Case in bench
[Uncaptioned image] [Uncaptioned image]
User How can I score? How can I draw a picture?
VGA-7b-v1 You can score the ‘dicatado’ button, which is located in the middle left of the page. You can click the blue ‘Start’ button in the center of the page to begin drawing a picture.
VGA-7b-v1* You should click the ‘dictado’ button located in the top left of the four options. You can start drawing by clicking the blue ’Start’ button
llava-mistral-v1.6-7b The click to score is the one that says "horrible" To draw a picture, you can use a drawing app on your phone or tablet.
Idefics2-8B To get the score, you should click on the "Score" button. To draw a picture, simply tap on the colors you want to use and then touch the screen where you would like to place the color
MiniCPM-V-2_5 To score, you click on the word .’horrible’ To draw a picture, you can use the drawing tools provided in the app interface.
Table 4: Case in the GUI-bench, note that VGA show the best comprehension of the image
6 Experiment
6.1 Baselines & Evaluation Metric
GUI comprehension Bench: To evaluate our model, we follow previous works Masry et al. (2023); Liu et al. (2023). Due to the lack of GUI comprehension bench, we sample 22 images from the Rico dataset (excluding training data). And based on these images, we collect 44 user questions which require truly understanding of GUI to response correctly. Inspired by the evaluation method of LLaVA-bench (in-the-wild) Liu et al. (2024b), we use ChatGPT to evaluate our model. We compare our model with five recent best performance LVLMs: the top non-open-source LVLMs GPT-4V and GPT-4o, the two best LVLMs based on Mistral-7b-instruction-v2 (llava-v1.6-mistral-7bLiu et al. (2024a) and idefics2-8bLaurençon et al. (2024)), and MiniCPM-llama3-V-2.5Hu et al. (2024), the leading open-source model on the LVLM leaderboard.
Fine-tuning Method Evaluation: To evaluate our two-stage fine-tuning approach (Foundation and Advanced Comprehension method), we track loss convergence over the training period. We compare loss convergence during the advanced task training between models pre-trained with the foundation task and which without it. We also compare previous models’ loss convergence with which trained on mixed foundation task and advanced task data. Their learning rates during training are all set to 2e-5, batch sizes are 16.
6.2 Main Result
GUI Comprehension: As shown in Table 3, VGA-7b-v1 has shown promising results, achieving the best performance across three separate GPT evaluations. Our model attains a score of 90.83, which is relatively 41% better than the base model llava-v1.6-mistral-7b, 54% better than idefics2-8b and 29% better than MiniCPM-llama3-V-2.5. Furthermore, VGA-7b-v1 also outperforms GPT-4o and GPT-4V. Overall, these results establish VGA as the SOTA model for GUI comprehension.
Resource Efficiency: We also evaluate our model’s performance with low-resolution input. As shown in Table 3, VGA-7b-v1* (336x336) still outperforms llava-v1.6-mistral-7b and idefics-8b, even though the number of image tokens for VGA-7b-v1* is only one-fourth of those in llava-v1.6-mistral-7b and idefics-8b. This significantly improves the inference speed, computational cost and memory usage. Our research indicates that reducing pixel count primarily affects detail recognition accuracy, thereby impacting response accuracy. Nevertheless, the model retains a strong understanding of GUI images, as demonstrated in Table 4.
Hallucination Analysis: Compared with existing models, our model demonstrates a precise understanding of GUI. As shown in Table 2 and Table 7, our model responds to questions based on the input image content and accurately describe GUI components using details such as relative position, color, and shape. By focusing more on image content, our model certainly reduces the likelihood of generating hallucinated responses.
FAC method analysis: As shown in Figure 2, the orange and purple curves represent the original loss values for models with and without the foundation task stage, respectively. The red and blue curves, fitted with a degree-3 polynomial, illustrate the loss trends. Model trained with foundation task exhibit more stable convergence and lower loss value. Figure 3 shows that mixing datasets during training leads to unstable convergence. In Appendix A.3, we discuses the influence of the foundation stage on model convergence. In Appendix A.4, we conducted ablation studies to evaluate the individual contributions of the two stages to overall model performance. In Appendix A.5, we analyzed the contribution of image tokens when generating answers to demonstrate how FAC enhances the model’s attention to image content.
7 Conclusion
We introduce VGA, a fine-tuned model for GUI comprehension, leveraging a custom dataset and novel fine-tuning method. The dataset, constructed using a Reference Method, focuses on image-centric responses. The training can be divided into two stage: foundational skill alignment and human intent alignment. The foundation stage emphasizes learning the relationship between images and responses, while the advanced stage focuses on intent following. Ablation studies confirm the efficacy of our training method. We believe that our training data and method will serve as valuable resources for future research.
Limitations
Despite the promising results, our work has some limitations that need to be addressed in future research. Firstly, our dataset contains some noisy data, which may affect the overall performance of the models. We plan to clean this data in future iterations to improve the quality and accuracy of the dataset. Secondly, the accuracy of responses is somewhat limited by the capability of the base model we used. Although our constructed dataset and proposed methodologies have demonstrated their effectiveness, there are still instances where the answers may be inaccurate. Future work will involve experimenting with more advanced base models to further validate and enhance the effectiveness of our approach. Due to the commercial licensing restrictions of the LLaMA model, we were unable to fine-tune based on the LLaMA model. At the time of our work, mistral-7b-instruction-v3 had not yet been released. Therefore, we utilized llava-v1.6-mistral-7b. In our research, it is one of the most advanced open-source and commercially viable models available. Thirdly, our dataset is constructed based on the open-source RICO dataset, which contains GUI images with older design styles. In future dataset construction, we aim to integrate GUI interfaces from current popular apps to ensure our models are up-to-date with modern design trends and practices. By addressing these limitations, we aim to refine our approach and further improve the effectiveness and reliability of our models in comprehending and interacting with complex GUI interfaces.
Ethics Statement
During the development process, we ensured compliance with the terms and conditions of the various models and datasets we utilized. The foundation model we employed, llava-v1.6-mistral-7b, is based on the mistral-7b-instruction-v2 and CLIP models. Both llava-v1.6-mistral-7b and mistral-7b-instruction-v2 are licensed under the Apache 2.0 License, which grants open-source and commercial-use permissions from both llava and mistral AI. Our dataset was constructed based on the RICO dataset. The University of Illinois provides a permissive license for the RICO dataset under the condition that users comply with their terms, allowing for open-source and commercial use. Due to the generative nature of our models, there is a potential risk that they may be misused to generate factually incorrect responses that could misinform the public. Additionally, we cannot guarantee that our models will not produce text containing hate speech or harmful content.
Acknowledgements
The authors would like to thank the anonymous reviewers for their helpful comments. This research was supported by Bytedance/TikTok, we would like to extend our gratitude to Bytedance/TikTok for their assistance and support.
Appendix A Appendices
A.1 Existing Method For Data Generation
To effectively train an LVLM, an extensive amount of data with high-quality is required Zhu et al. (2023). Due to the huge labor consumption to create a large high-quality dataset, many research initiatives have explored methods for automatic generation of datasets Gilardi et al. (2023)Chiang et al. (2023)Peng et al. (2023).
Current mainstream method focus on distilling knowledge from advanced large language models and large visual-language models, which are pre-trained on vast amounts of data, often exceeding terabyte (TB) scales, to capture a wide variety of patterns in language and visuals Xu et al. (2023)Chiang et al. (2023)Maeng et al. (2017)Chen et al. (2024). Therefore, they are often used as auxiliary tools to generate high-quality data. For instance, ShareGPT4v Chen et al. (2023) distill the knowledge from GPT-4V to generate data for training, ensuring the image and text pairs are of high-quality and diversity. LLaVA Liu et al. (2023) turns pictures into text descriptions and then uses GPT-4 OpenAI (2024) to generate questions and answers based on text descriptions. Both of them get a state-of-the-art performance.
A.2 Hyperparameter Analysis
Our fine-tuning process is based on the llava-v1.6-mistral-7b model. In this section, we examine the influence of different learning rates and batch sizes on model convergence across two stages.
Figures 14 and Figure 14 display the loss convergence during foundation task training at various learning rates, both of the batch sizes are 16. It is evident that a learning rate of 2e-6 achieves lower loss values and more stable learning, with significantly fewer fluctuations.
Figures 14 and Figure 14 show the loss convergence during advanced task training of models pre-trained with different learning rates on foundation task. The results indicate that model pre-trained with a learning rate of 2e-6 on foundation task results in markedly lower loss and reduced fluctuations during advanced task training.
Comparing Figure 14 and Figure 14, which both pre-trained with learning rate of 2e-6 on foundation task, and with the different batch size (32 and 16). This comparison shows that increasing the batch size significantly improves convergence stability and reduces fluctuations during advanced task, avoids over-fit in intent following.
In our experiments, the model trained under the conditions depicted in Figure 14 surpasses others. This model achieves higher accuracy in GUI element recognition compared to the which pre-trained on foundation task with a learning rate of 2e-5. and show more accuracy in human intent following compared with model trained with smaller batch size during advanced task training. This further underscores the importance of foundation task training and highlights the distinct focuses of the two stages. Using a larger batch size in advanced helps the model capture the nuances of user intent, thereby preventing over-fitting. Using a small learning rate in foundation task training helps the model more accurately map images to their corresponding responses.
A.3 Influence of Foundation Task
As shown in Figure 14, the performance during the foundation training phase significantly impacts the results of advanced task training. As shown in Figure 14 and Figure 14, if the model achieves stable convergence during the foundation task, it will also exhibit stable convergence during the advanced task. Furthermore, the final convergence value in the advanced task is influenced by the convergence value achieved during the foundation task. Using models with lower convergence values from the foundation task leads to smaller convergence values during the advanced task task training.
A.4 Ablation Study
We conduct ablation studies compare among base model (llava-v1.6-mistral-7b) and models fine-tuned by FAC method (VGA-7b-v1)and solely on advanced task. VGA-7b-v1 is fine-tuned based on model which pre-trained on foundation task, VGA-7b-stage2 is fine-tuned based on llava-v1.6-mistral-7b. As shown in Figure 7, we compare their loss convergence during advanced task and result shows that the two models exhibit highly consistent loss trends, but VGA-7b-v1 demonstrates lower loss values. This indicates that our foundation task training approach, which involves freezing the question and response formats, does not degrade the model’s intent-following performance. On the contrary, this training method allows the model to focus more on image information, leading to lower loss values during the advanced task training while maintaining similar trends.
Figure 4: VGA-7b-stage2
Figure 5: VGA-7b-v1 (with FAC method)
Figure 6: Loss convergence comparing between VGA-7b-v1 and VGA-7b-stage2
Figure 7:
We present some case in Table 9, we evaluate this three models on real-world tasks. Compared to the baseline model llava-v1.6-mistral-7b, the responses of VGA-7b-stage2 incorporate the answer style of the reference method. Compared to VGA-7b-v1, the models that not trained on foundation task shows limitations in GUI recognition accuracy, particularly in positional accuracy. This highlights the functionality and effectiveness of the two-phase training approach.
A.5 Attention Analysis
To verify that our method increases the model’s focus on extracting information from images when generating responses, we analyzed the attention values between image tokens and answer tokens. Specifically, we calculated the average attention contribution of each image token for every answer token during the generation process.. Additionally, we computed the total attention value of image tokens for each answer token during its generation.
As shown in Figure 23, we compared our model VGA-7b-v1 with the baseline model llava-v1.6-mistral-7b using the same images and questions as input, and we recorded the differences in attention value. The results show that our model has significantly higher attention values to image tokens when generating responses compared to the llava model. This indicates that our model is more capable of capturing the content of the images.
Experiment Data Size Training Model Learning Rate Batch Size GPUs Time
Foundation Task Training
Foundation Task 22.3k Connector& LLM 2e-6 16 1xA100 80G 16h
Advanced Comprehension Task Training
Advanced Task 41.4k Connector& LLM 2e-6 32 1xA100 80G 16h
Table 5: Hyperparameter in Training
Figure 8:
Figure 9:
Figure 10:
Figure 11:
Figure 12: VGA-7b-v1
Figure 13:
Figure 14: Train hyperparameter analysis
Figure 15: llava-v1.6-mistral-7b
Figure 16: VGA-7b-v1
Figure 17: llava-v1.6-mistral-7b
Figure 18: VGA-7b-v1
Figure 19: llava-v1.6-mistral-7b
Figure 20: VGA-7b-v1
Figure 21: llava-v1.6-mistral-7b
Figure 22: VGA-7b-v1
Figure 23:
Case in bench
[Uncaptioned image] [Uncaptioned image]
User What the day and date of the weekend before I selected in the middle of the page? How can I find my Parish?
VGA-7b-v1 It is SAT, Apr 9, 2016 You can push "SEARCH" button in the middle of the page with a blue text.
llava-mistral-v1.6-7b The image shows a calendar for April 2017, and you have selected the date April 10, which is a Monday. The weekend before April 10 is Saturday, April 8, and Sunday, April 9. You can find your Parish by customizing this app with relevant message, events, Mass times and more.
Idefics2-8B The day and date of the weekend before I selected are April 15th and 16th. Year 2016 Missed. To find your parish, you can use the search bar provided on the app. This will allow you to input the name of your parish or search for it using other relevant information.
MiniCPM-V-2_5 The day and date of the weekend before April 10, which is a Monday, would be Sunday, April 9. Year 2016 Missed. To find your Parish using the app, you can customize it with relevant messages, events, and Mass times by selecting a parish from the list provided.
Table 6: Case in the GUI-benchmark
Case in bench
[Uncaptioned image] [Uncaptioned image]
User How can I start Drawing? Am I use current location right now?
VGA-7b-v1 You can click the orange "Draw" button in the bottom left corner of the page. Yes, you are using you current location. The location icon is in the first line named "Use Current Location".
llava-mistral-v1.6-7b To start drawing, you can use the Draw app on your device, which is available in the app store. Yes, you are using your current location right now.
Idefics2-8B To start drawing, you can either tap on the "Draw" button or select one of the three options provided (Coloring, My Work, or Start Now). Yes, I am using your current location to provide information about nearby places of interest.
MiniCPM-V-2_5 You can click the "Draw" button. No, you are not currently using your current location.
Table 7: Case in the GUI-benchmark
Case in bench
[Uncaptioned image] [Uncaptioned image]
User How can I remove the vehicle? Where can I find the work I drew?
VGA-7b-v1 You can remove the vehicle by clicking on the ’Delete Vehicle’ option in the dropdown menu at the top right of the page. It’s the third option from the top. You can find the work you’ve drawn by clicking on the orange square icon labeled ’My Work’ at the bottom right of the page.
GPT-4o To remove the vehicle, tap on "Delete Vehicle" from the menu shown. You can find the paintings you drew by clicking on the "My Work" button.
GPT-4V You can remove the vehicle by selecting the "Delete Vehicle" option in the menu shown on the app. You can find the paintings you drew in the "My Work" section of the app shown in the image.
MiniCPM-V-2_5 To remove a vehicle from the application, you would typically select the ’Delete Vehicle’ option from the menu. The "My Work" icon suggests a section where users can view their past drawings.
Table 8: Case in the GUI-benchmark
Ablation Study
[Uncaptioned image] [Uncaptioned image]
User Who will play against each other on Sun, Aug 14? where is the diocese?
VGA-7b-v1 The Cowboys and the Rams will play against each other on sun, aug 14. The Diocese icon is in the bottom row, second from the left. It is a white icon with a video camera symbol.
VGA-7b-stage2 The Cowboys will play against the Dolphins on Sun, Aug 14.. The Diocese is in the second row, second column.
llava-v1.6-mistral-7b Cowboys and dolphins. The diocese icon is located in the bottom left corner of the app.
Table 9: Ablation study of the model trained with foundation task and without. llava-v1.6-mistral-7b is the base model, VGA-7B-stage2 is the model trained solely on advanced task.
Real world
[Uncaptioned image] [Uncaptioned image]
User Please open the weather app, click? What is the weather on Sat?
VGA-7b-v1 <0.12, 0.28> The weather in Sat is sunny with a temperature of 22°C.
User Has anyone called me? What is the highest temperature today?
VGA-7b-v1 Yes, there is a call icon with a red notification badge on it. The highest temperature today is 30°C. You can find it in the bottom section of the page, under the ‘10-DAY FORECAST’ section.
User Could you help me to play some song? what is the temperature in 19:00?
VGA-7b-v1 To play a song, tap on the icon that looks like a music note. It’s located in the bottom row, second from the left. The temperature in 19:00 is 25°C.
User What is date of the week today? What is the weather like now?
VGA-7b-v1 Today’s date of the week is Thursday. The current weather is 25°C and cloudy.
User Open the camera. How will the weather change next?
VGA-7b-v1 To open the camera, tap on the screen at <0.34, 0.7>. The weather will continue for the rest of the day. Wind gusts are up to 7 m/s.
User What is the setting bounds? How about the air quality?
VGA-7b-v1 The setting bounds are <0.762, 0.566, 0.922, 0.692>. The air quality is currently labeled as ’Excellent’ with a value of 43.
Table 10: Case in the real world